The air in a 6.00 L tank has a pressure of 1.50 atm . What is the final pressure, in atmospheres, when the air is placed in tanks that have the following volumes, if there is no change in temperature and amount of gas?
1
Expert's answer
2020-04-02T09:00:17-0400

At constant temperature,By Boyle's law,

"P_1V_1=P_2V_2"

And in this case,"P_1=1.5\ atm,V_1=6L"

When "V_2=1L,P_2=\frac{P_1V_1}{V_2}=\frac{1.5\times6}{1}=9\ atm"

When "V_2=2500\ ml=2.5L,P_2=\frac{P_1V_1}{V_2}=\frac{1.5\times6}{2.5}=3.6\ atm"